SCI Question 3

SOLVE FOR COST OF GOODS SOLD

 

Gross Profit = 60% of net sales

Sales Discount = 120,000

Gross Sales= 900% of Sales Discounts


1
Expert's answer
2021-09-27T18:46:04-0400

Solution:

Cost of goods sold = Net sales – Gross profit

Derive Gross profit and Net sales:

Gross sales = 900"\%" of Sales Discounts

Sales Discount = 120,000

Gross sales = 900"\% \times" 120,000 = 1,080,000

Net sales = Gross sales – Sales Discounts = 1,080,000 – 120,000 = 960,000

Gross Profit = 60"\%" of net sales = 60"\% \times"960,000 = 576,000

Cost of goods sold = Net sales – Gross Profit = 960,000 – 576,000 = 384,000

 

Cost of goods sold = 384,000